Please help me with this algebra problem. It states:
If a probasketball plaer has a vertical leap of about 25 inches, what is his hang time? Use the hang-time function V=48T^2.
Please help solve. Thanks.

Answer by nerdybill(7384)

Your problem gives you:
V=48T^2
where
V is the "vertical leap"
T is the "hang time"
.
Now, they want to know the "hang time" if the "vertical leap" is 25 inches. So, plug in the given value (V) and solve for T:
V=48T^2
25=48T^2
25/48 = T^2
0.521 = T^2
sqrt(0.521) = T
0.722 seconds = T